OpenStudy (anonymous):

The perimeter of a square is 160 feet. What is the area of the square

OpenStudy (campbell_st):

side length is 160/4 Area = (160/4)^2

OpenStudy (unklerhaukus):

\[Perimeter_{square} =4s\]\[Area_{square} =s^2\] \[P=4s=160ft~~~\rightarrow s=40ft\]\[A=(40ft)^2=1600ft^2\]

OpenStudy (unklerhaukus):